A sneaker shop pays its wholesaler $40 for a certian athletic shoe, and then sells it to a customer for $75. What is the % markup?
Answer by Alan3354(69443)

A sneaker shop pays its wholesaler $40 for a certian athletic shoe, and then sells it to a customer for $75. What is the % markup?
----------
The markup is $35 (75 - 40)
35/40 = 7/8 = 0.875
= 87.5%